OpenStudy (anonymous):

What is the area of a circle with a diameter of 4?

OpenStudy (anonymous):

if diameter is given then 1st of all you find the radius and the diameter will be divided by 2 so the radius will be 2 and then apply the farmula of circumference 2pi r 2*22/7*2=---------- then put the farmula of area that is pi*r squ.=22/7*4=---------- after calculating you will find the answer that is 12.57cm square.

OpenStudy (disco619):

First we would need to find "radius" of circle. For that we would divide diameter by 2 as diameter is the length from one side to the other side of circle passing through the circle whereas radius is from one side to center.  So, \[Diameter = 4\] \[Radius = \frac{ diameter }{ 2 }\] \[Radius = \frac{ 4 }{ 2 }\] \[Radius = 2\] Now we would use the Formula of Area of Circle which is: \[\pi r^{2}\] \[\pi \times radius ^{2} \] \[\pi \times 2^{2}\] \[\pi \times 4\] \[= 12.5663706144\] Now it depends if you have to round the answer or not. If you have to round it then it would be: \[= 13\] And it also depends what is the diameter in. In centimeter, meter, feet or what? If the diameter is in cm and the answer has to be rounded off, then you would write as this: \[= 13 cm ^{2} \]

OpenStudy (anonymous):

THANK YOU ! @Disco619

OpenStudy (disco619):

Any Time... :)